OpenStudy (maddie82502):

A vase contains 12 tulips and 20 roses. What is the ratio of roses to the total number of flowers in the vase? A. 3 to 8 B. 3 to 5 C. 5 to 8 D. 5 to 3

OpenStudy (peach3113):

Imagine 12/20. Reduce the fraction.

OpenStudy (peach3113):

It can be reduced by 4. 12÷4=? 20÷4=?

OpenStudy (peach3113):

12÷4=3 20÷4=5

OpenStudy (maddie82502):

3 to 5

OpenStudy (peach3113):

Yeppers! :D

OpenStudy (peach3113):

Any more questions? :D